### v1

@@ 3.1 Approximation of SSS matrices
(5) $\operatorname{\mathbb{E}}\Bigl{[}\|\bm{A}-(\bm{U}\bm{X}\bm{V}^{\mathsf{T}}+\bm{D})\|_{\mathsf{F}}^{2}\Bigr{]}\leq(\Gamma_{\textup{r}}+\Gamma_{\textup{c}})(1+\Gamma_{\textup{d}})\cdot\inf\limits_{\bm{B}\in\operatorname{SSS}(\ell,k)}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}.$
Proof 3.3.
@@
$\|\bm{A}-(\widetilde{\bm{U}}\widetilde{\bm{X}}\widetilde{\bm{V}}^{\mathsf{T}}+\widetilde{\bm{D}})\|_{\mathsf{F}}^{2}=\inf\limits_{\bm{B}\in\operatorname{SSS}(\ell,k)}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}=:\text{OPT}^{2}.$
@@
We begin by bounding $R$. Define $\widetilde{\bm{X}}_{i,:}=\begin{bmatrix}\widetilde{\bm{X}}_{i,1}&\widetilde{\bm{X}}_{i,2}&\cdots&\widetilde{\bm{X}}_{i,2^{L}}\end{bmatrix}$ Using our assumption on $\bm{U}_{i}$ and that $[\![r_{i}(\bm{A})]\!]_{k}$ is the best rank-$k$ approximation to $r_{i}(\bm{A})$, we have $\operatorname{\mathbb{E}}\Bigl{[}\|r_{i}(\bm{A})-\bm{U}_{i}\bm{U}_{i}^{\mathsf{T}}r_{i}(\bm{A})\|_{\mathsf{F}}^{2}\Big{]}\leq\Gamma_{\textup{r}}\cdot\|r_{i}(\bm{A})-[\![r_{i}(\bm{A})]\!]_{k}\|_{\mathsf{F}}^{2}$. Then, since $\widetilde{\bm{U}}_{i}\widetilde{\bm{X}}_{i,:}\widetilde{\bm{V}}^{\mathsf{T}}$ is some rank-$k$ approximation to $r_{i}(\bm{A})$, and recalling that $\widetilde{\bm{X}}_{i,i}=\bm{0}$,
$R$ $\leq\Gamma_{\textup{r}}(1+\Gamma_{\textup{d}})\cdot\sum_{i=1}^{2^{L}}\|r_{i}(\bm{A})-[\![r_{i}(\bm{A})]\!]_{k}\|_{\mathsf{F}}^{2}$
@@
Remark 3.4.
The proof of Theorem 3.2 can be used to obtain an absolute error bound, similar to those found in the literature [40]. Suppose that we have explicit access to the entries of the matrix $\bm{A}$ so that we can set $\Gamma_{\textup{d}}=0$ and that $k$ is so that for each $i$, we have $\|r_{i}(\bm{A})-\bm{U}_{i}\bm{U}_{i}^{\mathsf{T}}r_{i}(\bm{A})\|_{\mathsf{F}}\leq\varepsilon$ and $\|c_{i}(\bm{A})-c_{i}(\bm{A})\bm{V}_{i}\bm{V}_{i}^{\mathsf{T}}\|_{\mathsf{F}}\leq\varepsilon$. Then, following the proof of Theorem 3.2, we have $\|\bm{A}-(\bm{U}\bm{X}\bm{V}^{\mathsf{T}}+\bm{D})\|_{\mathsf{F}}\leq\sqrt{2^{\ell+1}}\varepsilon$.
@@ 3.2 Bounding the HSS optimum at different levels
$\inf\limits_{\bm{C}\in\operatorname{HSS}(\ell-1,k)}\|\bm{U}^{\mathsf{T}}(\bm{A}-\bm{D})\bm{V}-\bm{C}\|_{\mathsf{F}}\leq\inf\limits_{\bm{B}\in\operatorname{HSS}(\ell,k)}\|\bm{A}-\bm{B}\|_{\mathsf{F}}.$
To prove Lemma 3.5, we first show that certain affine transformations of HSS matrices are also HSS. A similar result is found in [40, Lemma 4.6] and, for completeness, we include the proof in Subsection A.1.
@@
Fix some $\varepsilon>0$ and let $\widetilde{\bm{B}}\in\operatorname{HSS}(\ell,k)$ be a near-optimal HSS approximation to $\bm{A}$ so that $\|\bm{A}-\widetilde{\bm{B}}\|_{\mathsf{F}}^{2}\leq(1+\varepsilon)\inf\limits_{\bm{B}\in\operatorname{HSS}(\ell,k)}\|\bm{A}-\bm{B}\|_{\mathsf{F}}$. [^5] Recall that we are not aware of any results showing that an optimal HSS approximation exists. However, a $(1+\epsilon)$-optimal approximation must exist for any $\epsilon>0$. Define $\widetilde{\bm{E}}=\bm{A}-\widetilde{\bm{B}}$. Then,
@@
$\inf\limits_{\bm{C}\in\operatorname{HSS}(\ell-1,k)}\|\bm{U}^{\mathsf{T}}(\bm{A}-\bm{D})\bm{V}-\bm{C}\|_{\mathsf{F}}$ $\leq\|\bm{U}^{\mathsf{T}}(\bm{A}-\bm{D})\bm{V}-\bm{U}^{\mathsf{T}}(\widetilde{\bm{B}}-\bm{D})\bm{V}\|_{\mathsf{F}}$
@@
$\leq\|\widetilde{\bm{E}}\|_{\mathsf{F}}=(1+\varepsilon)\inf\limits_{\bm{B}\in\operatorname{HSS}(\ell,k)}\|\bm{A}-\bm{B}\|_{\mathsf{F}}.$
Letting $\varepsilon\to 0$ yields the desired result.
@@ 3.3 Analysis of Algorithm 1
Now assume, for the sake of induction, that the theorem holds for any approximation in $\operatorname{HSS}(L-1,k)$ that is outputted by Algorithm 1. That is, for any such approximation, the bound Eq. 3 holds with multiplicative factor $(\Gamma_{r}+\Gamma_{c})(1+\Gamma_{d})(L-1)$. Write $\bm{B}$ in its telescoping factorization Equation 2 and $\bm{X}=(\bm{U}^{(L)})^{\mathsf{T}}(\bm{A}-\bm{D})\bm{V}^{(L)}$. Then,
@@
$\|\bm{A}-\bm{D}^{(L)}-\bm{U}^{(L)}\bm{X}(\bm{V}^{(L)})^{\mathsf{T}}\|_{\mathsf{F}}^{2}$ $\leq(1+\Gamma_{\textup{d}})(\Gamma_{\textup{c}}+\Gamma_{\textup{r}})\inf\limits_{\bm{C}\in\operatorname{SSS}(L,k)}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2}$
(12) $\leq(1+\Gamma_{\textup{d}})(\Gamma_{\textup{c}}+\Gamma_{\textup{r}})\inf\limits_{\bm{C}\in\operatorname{HSS}(L,k)}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2}.$
Note that $\bm{B}^{(L)}\in\operatorname{HSS}(L-1,k)$ is the output of Algorithm 1 when we run it on $\bm{A}$. Therefore, by our inductive assumption and Lemma 3.5 we have
$\|\bm{X}-\bm{B}^{(L)}\|_{\mathsf{F}}^{2}$ $\leq(1+\Gamma_{\textup{d}})(\Gamma_{\textup{c}}+\Gamma_{\textup{r}})(L-1)\inf\limits_{\bm{C}\in\operatorname{HSS}(L-1,k)}\|\bm{X}-\bm{C}\|_{\mathsf{F}}^{2}$
(13) $\leq(1+\Gamma_{\textup{d}})(\Gamma_{\textup{c}}+\Gamma_{\textup{r}})(L-1)\inf\limits_{\bm{C}\in\operatorname{HSS}(L,k)}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2}.$
@@
Suppose that we have explicit access to the entries of the matrix $\bm{A}$ so that we can set $\Gamma_{\textup{d}}=0$. Then, suppose that for each block row and column, we have
@@
$\|c_{i}(\bm{A}^{(\ell+1)})-c_{i}(\bm{A}^{(\ell+1)})\bm{V}_{i}^{(\ell)}\bm{V}_{i}^{(\ell)\mathsf{T}}\|_{\mathsf{F}}$ $\leq\varepsilon.$
Then, by the proof of Theorem 3.1 and Remark 3.4 we have that the output of Algorithm 1 satisfies
$\|\bm{A}-\bm{B}\|_{\mathsf{F}}\leq\bigg{(}\sum\limits_{\ell=1}^{L}2^{\ell+1}\bigg{)}^{1/2}\cdot\varepsilon\leq 2^{(L+2)/2}\cdot\varepsilon=\sqrt{2N/k}\cdot\varepsilon,$
since $N=2^{L+1}k$. In particular, if we truncate all the singular values of each HSS-block row and HSS-block columns less than $\epsilon$ we have $\varepsilon\leq\sqrt{2k}\epsilon$ and we therefore have $\|\bm{A}-\bm{B}\|_{\mathsf{F}}\leq 2\sqrt{N}\epsilon$. This provides a similar bound to [40, Theorem 4.7], which gives a $O(\sqrt{N}\epsilon)$ bound in the spectral norm.
@@ 3.4 Lower bound for Algorithm 1
$\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}\geq(2-\varepsilon)\inf\limits_{\bm{C}\in\operatorname{HSS}(L,k)}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2}.$
The proof, which is given in Subsection A.2, is fully constructive, and in fact provides a low-bound for the greedy SSS approximation algorithm used at each iteration of Algorithm 1. In general, we are unaware of any polynomial-time algorithm which can compute a $(1+\varepsilon)$ factor HSS or SSS approximation. In fact, it is even an open question whether an any $O(1)$ approximation factor is possible for HSS matrices.
@@ 4 A matrix-vector query algorithm
$\mathbb{E}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}\leq\Gamma\cdot\inf\limits_{\bm{C}\in\operatorname{HSS}(L,k)}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2},\quad\text{where}\quad\Gamma=O(L).$
This remainder of this section is structured as follows. In Subsection 4.1, we review the technique of block nullification, which allows us to efficiently simulate matvecs with the block-rows and block-columns of $\bm{A}$ using matrix-vector products with the full matrix. In Subsection 4.2, we establish explicit error bounds (Theorem 4.2) for the PCPS-based low-rank approximation technique that is used in Algorithm 2 to obtain subspaces for the HSS block-rows and block-columns. We then analyze an approach for approximating the diagonal blocks at each level in Subsection 4.3. Finally, in Subsection 4.4, we prove our main result, Theorem 4.1, by showing how the error bound for these primitives allows us to instantiate Theorem 3.1.

### v2

@@ 3.1 Approximation of SSS matrices
(5) $\operatorname{\mathbb{E}}\Bigl{[}\|\bm{A}-(\bm{U}\bm{X}\bm{V}^{\mathsf{T}}+\bm{D})\|_{\mathsf{F}}^{2}\Bigr{]}\leq(\Gamma_{\textup{r}}+\Gamma_{\textup{c}})(1+\Gamma_{\textup{d}})\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{B}\in\operatorname{SSS}(\ell,k)}}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}.$
Proof 3.4.
@@
$\|\bm{A}-(\widetilde{\bm{U}}\widetilde{\bm{X}}\widetilde{\bm{V}}^{\mathsf{T}}+\widetilde{\bm{D}})\|_{\mathsf{F}}^{2}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{B}\in\operatorname{SSS}(\ell,k)}}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}=:\text{OPT}^{2}.$
@@
We begin by bounding $R$. Define $\widetilde{\bm{X}}_{i,:}=\begin{bmatrix}\widetilde{\bm{X}}_{i,1}&\widetilde{\bm{X}}_{i,2}&\cdots&\widetilde{\bm{X}}_{i,2^{L}}\end{bmatrix}$ Using our assumption on $\bm{U}_{i}$ and that $r_{i}(\bm{A})]\!]_{k}$ is the best rank-$k$ approximation to $r_{i}(\bm{A})$, we have $\operatorname{\mathbb{E}}\Bigl{[}\|r_{i}(\bm{A})-\bm{U}_{i}\bm{U}_{i}^{\mathsf{T}}r_{i}(\bm{A})\|_{\mathsf{F}}^{2}\Big{]}\leq\Gamma_{\textup{r}}\cdot\|r_{i}(\bm{A})-[\![r_{i}(\bm{A})]\!]_{k}\|_{\mathsf{F}}^{2}$. Then, since $\widetilde{\bm{U}}_{i}\widetilde{\bm{X}}_{i,:}\widetilde{\bm{V}}^{\mathsf{T}}$ is some rank-$k$ approximation to $r_{i}(\bm{A})$, and recalling that $\widetilde{\bm{X}}_{i,i}=\bm{0}$,
$R$ $\leq\Gamma_{\textup{r}}(1+\Gamma_{\textup{d}})\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sum_{i=1}^{2^{\ell}}\|r_{i}(\bm{A})-[\![r_{i}(\bm{A})]\!]_{k}\|_{\mathsf{F}}^{2}}$
@@ 3.2 Bounding the HSS optimum at different levels
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{C}\in\operatorname{HSS}(\ell-1,k)}}\|\bm{U}^{\mathsf{T}}(\bm{A}-\bm{D})\bm{V}-\bm{C}\|_{\mathsf{F}}\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{B}\in\operatorname{HSS}(\ell,k)}}\|\bm{A}-\bm{B}\|_{\mathsf{F}}.$
To prove Lemma 3.5, we first show that certain affine transformations of HSS matrices are also HSS. A similar result is found in [42, Lemma 4.6] and, for completeness, we include the proof in Subsection A.1.
@@
Let $\widetilde{\bm{B}}\in\operatorname{HSS}(\ell,k)$ be an optimal HSS approximation to $\bm{A}$ so that $\|\bm{A}-\widetilde{\bm{B}}\|_{\mathsf{F}}^{2}=\min\limits_{\bm{B}\in\operatorname{HSS}(\ell,k)}\|\bm{A}-\bm{B}\|_{\mathsf{F}}$. [^6] Recall that the result in Appendix D guarantees the existence of an optimal approximation. Define $\widetilde{\bm{E}}=\bm{A}-\widetilde{\bm{B}}$. Then,
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{C}\in\operatorname{HSS}(\ell-1,k)}}\|\bm{U}^{\mathsf{T}}(\bm{A}-\bm{D})\bm{V}-\bm{C}\|_{\mathsf{F}}$ $\leq\|\bm{U}^{\mathsf{T}}(\bm{A}-\bm{D})\bm{V}-\bm{U}^{\mathsf{T}}(\widetilde{\bm{B}}-\bm{D})\bm{V}\|_{\mathsf{F}}$
@@
$\leq\|\widetilde{\bm{E}}\|_{\mathsf{F}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{B}\in\operatorname{HSS}(\ell,k)}\|\bm{A}-\bm{B}\|_{\mathsf{F}},}$
which yields the desired result.
@@ 3.3 Analysis of Algorithm 1
Now assume, for the sake of induction, that the theorem holds for any approximation in $\operatorname{HSS}(L-1,k)$ that is outputted by Algorithm 1. That is, for any such approximation, the bound Eq. 3 holds with multiplicative factor $(\Gamma_{r}+\Gamma_{c})(1+\Gamma_{d})(L-1)$. Write $\bm{B}$ in its telescoping factorization Equation 2 and $\bm{X}=(\bm{U}^{(L)})^{\mathsf{T}}(\bm{A}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{D}^{(L)}})\bm{V}^{(L)}$. Then,
@@
$\|\bm{A}-\bm{D}^{(L)}-\bm{U}^{(L)}\bm{X}(\bm{V}^{(L)})^{\mathsf{T}}\|_{\mathsf{F}}^{2}$ $\leq(1+\Gamma_{\textup{d}})(\Gamma_{\textup{c}}+\Gamma_{\textup{r}}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{C}\in\operatorname{SSS}(L,k)}}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2}$
(12) $\leq(1+\Gamma_{\textup{d}})(\Gamma_{\textup{c}}+\Gamma_{\textup{r}}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{C}\in\operatorname{HSS}(L,k)}}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2}.$
Note that $\bm{B}^{(L)}\in\operatorname{HSS}(L-1,k)$ is the output of Algorithm 1 when we run it on $\bm{X}$. Therefore, by our inductive assumption and Lemma 3.5 we have
$\|\bm{X}-\bm{B}^{(L)}\|_{\mathsf{F}}^{2}$ $\leq(1+\Gamma_{\textup{d}})(\Gamma_{\textup{c}}+\Gamma_{\textup{r}})(L-1){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{C}\in\operatorname{HSS}(L-1,k)}}\|\bm{X}-\bm{C}\|_{\mathsf{F}}^{2}$
(13) $\leq(1+\Gamma_{\textup{d}})(\Gamma_{\textup{c}}+\Gamma_{\textup{r}})(L-1){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{C}\in\operatorname{HSS}(L,k)}}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2}.$
@@
The proof of Theorem 3.1 can be modified to obtain an absolute error bound, similar to those found in the literature [42]. Suppose that we have explicit access to the entries of the matrix $\bm{A}$ so that we can set $\Gamma_{\textup{d}}=0$. Suppose further that we can compute orthonormal bases $\bm{U}_{i}^{(\ell)}$ and $\bm{V}_{i}^{(\ell)}$ so that
@@
$\|c_{i}(\bm{A}^{(\ell+1)})-c_{i}(\bm{A}^{(\ell+1)})\bm{V}_{i}^{(\ell)}\bm{V}_{i}^{(\ell)\mathsf{T}}\|_{\mathsf{F}}$ $\leq\varepsilon,$
for each $i$ and $\ell$. By the arguments of Theorem 3.3, we have
$\|\bm{A}^{(\ell+1)}-\bm{U}^{(\ell)}\bm{A}^{(\ell)}\left(\bm{V}^{(\ell)}\right)^{\mathsf{T}}-\bm{D}^{(\ell)}\|_{\mathsf{F}}^{2}\leq$ $\sum\limits_{i=1}^{2^{\ell}}\left[\|r_{i}(\bm{A}^{(\ell+1)})-\bm{U}_{i}^{(\ell)}\bm{U}_{i}^{(\ell)\mathsf{T}}r_{i}(\bm{A}^{(\ell+1)})\|_{\mathsf{F}}^{2}\right$.
$\hskip-30.00005pt+\left.\|c_{i}(\bm{A}^{(\ell+1)})-c_{i}(\bm{A}^{(\ell+1)})\bm{V}_{i}^{(\ell)}\bm{V}_{i}^{(\ell)\mathsf{T}}\|_{\mathsf{F}}^{2}\right]$
$\leq$ $2\sum\limits_{i=1}^{2^{\ell}}\varepsilon^{2}=2^{\ell+1}\varepsilon^{2}.$
Combining this with the proof of Theorem 3.1, the output of Algorithm 1 satisfies
$\|\bm{A}-\bm{B}\|_{\mathsf{F}}\leq\bigg{(}2^{L+1}+\sum\limits_{\ell=1}^{L-1}2^{\ell+1}\bigg{)}^{1/2}\cdot\varepsilon+\bigg{(}\sum\limits_{\ell=1}^{L}2^{\ell+1}\bigg{)}^{1/2}\cdot\varepsilon\leq 2^{(L+2)/2}\cdot\varepsilon=\sqrt{2N/k}\cdot\varepsilon,$
since $N=2^{L+1}k$. In particular, if we truncate all the singular values of each HSS-block row and HSS-block columns less than $\epsilon$, we have $\varepsilon\leq\sqrt{2k}\epsilon$ and we therefore have $\|\bm{A}-\bm{B}\|_{\mathsf{F}}\leq 2\sqrt{N}\epsilon$. This provides a similar bound to [42, Theorem 4.7], which gives a $O(\sqrt{N}\epsilon)$ bound in the spectral norm.
@@ 3.4 Lower bound for Algorithm 1
$\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}\geq(2-\varepsilon){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{C}\in\operatorname{HSS}(L,k)}}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2}.$
The proof, which is given in Subsection A.2, is fully constructive and in fact provides a lower bound for the greedy SSS approximation algorithm used at each iteration of Algorithm 1. In general, we are unaware of any polynomial-time algorithm which can compute a $(1+\varepsilon)$ factor HSS or SSS approximation. In fact, it is even an open question whether an any $O(1)$ approximation factor is possible for HSS matrices.
We note that our lower bound in Theorem 3.10 does not depend on $L$, while the error in our upper bound scales as $O(L)$. Addressing this gap and finding the correct dependence on $L$ (if there is any at all) is another interesting open problem.
@@ 4 A matrix-vector query algorithm
$\mathbb{E}\|\bm{A}-\bm{B}\|_{\mathsf{F}}^{2}\leq\Gamma\cdot{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\min\limits_{\bm{C}\in\operatorname{HSS}(L,k)}}\|\bm{A}-\bm{C}\|_{\mathsf{F}}^{2},\quad\text{where}\quad\Gamma=O(L).$
The remainder of this section is structured as follows. In Subsection 4.1, we review the technique of block nullification, which allows us to efficiently simulate matvecs with the block-rows and block-columns of $\bm{A}$ using matrix-vector products with the full matrix. In Subsection 4.2, we establish explicit error bounds (Theorem 4.2) for the PCPS-based low-rank approximation technique that is used in Algorithm 2 to obtain subspaces for the HSS block-rows and block-columns. We then analyze an approach for approximating the diagonal blocks at each level in Subsection 4.3. Finally, in Subsection 4.4, we prove our main result, Theorem 4.1, by showing how the error bound for these primitives allows us to instantiate Theorem 3.1.